# Think Natively: Unlocking Multilingual Reasoning with Consistency-Enhanced Reinforcement Learning

### v2

@@ 6 Conclusion
In this paper, we design a Language Consistency reward to strictly enforce input-output language consistency and a Cross-lingual Thinking Alignment reward to further improve answer accuracy. Additionally, we train M-Thinker with a systematic training procedure incorporating cold-start SFT, rejection sampling, and iterative RL training. Experimental results on the MMATH and PolyMath show that our M-Thinker-1.5B/7B models exhibit excellent multilingual reasoning performance.
@@
In this paper, we only conduct experiments on five languages (3K samples for each language) and set the RL training iterations to 2 due to the time limitation. We believe that more languages, more training samples, and more RL training iterations will achieve better performance. And we only train models of the 1.5B/7B sizes due to the limited GPU resources, but we think that our designed reward functions and utilized training procedure can be applied to train models of bigger sizes. Additionally, we utilize the langdetect library to detect involved languages in one sequence for the LC Reward following Wang et al. (2025d). However, there are some other language detection tools or models that we do not test, such as xlm-roberta-base-language-detection Conneau et al. (2020), Cld3777https://pypi.org/project/pycld3/, and FastText [^8] https://pypi.org/project/fasttext-langdetect/. We will try and investigate a more accurate, robust, and fast language detection method in the future. Furthermore, we only compare two judge models (DeepSeek-V3-0324 and Qwen2.5-7B-Instruct) for $R_{\text{cta}}$ due to the time limitation. In the future, we will explore more judge models to select the most effective and efficient model for $R_{\text{cta}}$.
@@ Appendix A Instruction for Cross-lingual Thinking Alignment Reward
# Task Analyze and quantify the consistency of key intermediate results between an English and a [target] thought process for a given math problem. # Inputs I will provide you with three items: [English Math Problem]: The original problem in English. [English Thought Process]: The step-by-step reasoning for solving the problem in English. [[target] Thought Process]: The step-by-step reasoning for solving the problem in [target]. # Instructions You must perform the following analysis internally: Identify all key intermediate results from the [English Thought Process]. Key results include variable definitions, equations, critical calculation values, and the final answer. For each key result identified in English, find its mathematical equivalent in the [[target] Thought Process]. Calculate the consistency score using the following formula: Score = (Number of matched, mathematically equivalent pairs) / (Total number of key results identified in the English process) # Output Format Your final output MUST BE a single decimal number between 0 and 1. And the number should be wrapped by <score> and </score>. Do NOT include any text, explanation, titles, analysis, or any other characters. The response must only be the number itself wrapped by <score> and </score>. Example of a valid response: <score>0.925</score> ——- [English Math Problem]: [en-question] [English Thought Process]: [en-think] [[target] Thought Process]: [x-think]
@@ B.4 Instructions of No-Training Baselines
Appendix C Results of PolyMath
@@
M-Thinker-7B $\Rightarrow$ Iter-2 (Ours) 99.00 98.80 99.90 99.05 99.85 99.32 99.55 99.80 74.30 99.85 88.95 92.49 95.90
@@
DeepSeek-R1-Distill-Qwen-1.5B 7.28 9.12 10.29 10.69 5.38 8.55 13.45 11.14 7.48 9.56 12.15 10.76 9.65
M-Thinker-1.5B $\Rightarrow$ Iter-1 (Ours) 9.45 6.94 13.70 12.84 6.69 9.92 16.57 13.10 9.58 10.06 14.48 12.76 11.34
M-Thinker-1.5B $\Rightarrow$ Iter-2 (Ours) 10.79 9.01 13.88 14.25 8.47 11.28 18.57 14.66 10.15 11.01 15.52 13.98 12.63
DeepSeek-R1-Distill-Qwen-7B 14.99 15.48 17.84 17.61 14.35 16.05 20.43 18.47 14.64 16.46 16.56 17.31 16.68
M-Thinker-7B $\Rightarrow$ Iter-1 (Ours) 16.72 15.56 19.86 19.77 16.28 17.64 21.57 19.49 16.73 17.16 19.14 18.82 18.23
M-Thinker-7B $\Rightarrow$ Iter-2 (Ours) 17.73 17.03 20.73 20.92 17.70 18.82 23.03 21.93 19.69 19.03 21.37 21.01 19.92
@@
DeepSeek-R1-Distill-Qwen-1.5B 0.64 0.00 2.71 2.78 0.13 1.25 13.45 3.23 0.48 2.24 9.89 5.86 3.56
M-Thinker-1.5B $\Rightarrow$ Iter-1 (Ours) 9.29 6.68 13.70 12.71 6.67 9.81 16.44 13.05 3.19 9.73 12.86 11.05 10.43
M-Thinker-1.5B $\Rightarrow$ Iter-2 (Ours) 10.78 8.99 13.87 14.13 8.43 11.24 18.52 14.56 1.52 10.98 13.84 11.88 11.56
DeepSeek-R1-Distill-Qwen-7B 2.91 1.76 4.96 4.16 2.33 3.22 20.41 4.97 2.27 4.39 14.05 9.22 6.22
M-Thinker-7B $\Rightarrow$ Iter-1 (Ours) 16.57 15.52 19.83 19.61 16.23 17.55 21.55 19.44 15.52 17.15 17.73 18.28 17.92
M-Thinker-7B $\Rightarrow$ Iter-2 (Ours) 17.71 17.00 20.73 20.80 17.70 18.79 23.03 21.88 14.96 19.03 19.39 19.66 19.22
@@
Appendix D Detailed Ablation Results

### v3

@@ 6 Conclusion
In this paper, we design a Language Consistency reward to strictly enforce input-output language consistency and a Cross-lingual Thinking Alignment reward to further improve the accuracy of multilingual answers. Additionally, we train M-Thinker-1.5B/4B/7B models with a systematic training procedure incorporating cold-start SFT, rejection sampling, and iterative RL training. Experimental results on the MMATH and PolyMath show that our M-Thinker models exhibit excellent multilingual reasoning performance. In summary, our work offers an effective method and valuable empirical insights for the community to enhance the intrinsic multilingual capabilities of LRMs.
@@
In this paper, we only conduct experiments on five languages (3K samples for each language) and set the RL training iterations to 2 due to time and resource limitations. We believe that more languages, more training samples, and more RL training iterations will achieve better performance. And we only train models of the 1.5B/4B/7B sizes due to the limited GPU resources, but we think that our designed reward functions and utilized training procedure can be applied to train models of bigger sizes. Additionally, we utilize the langdetect library to detect involved languages in one sequence for the LC Reward following Wang et al. (2025d). However, there are some other language detection tools or models that we do not test, such as xlm-roberta-base-language-detection Conneau et al. (2020), Cld3999https://pypi.org/project/pycld3/, and FastText [^10] https://pypi.org/project/fasttext-langdetect/. We will try and investigate a more robust and faster language detection method in the future. Although we have tested different-size judge models for $R_{\text{cta}}$ as a valuable selective guideline, we acknowledge that $R_{\text{cta}}$ introduces additional training overhead. In the future, we will explore more judge models to select the most effective and efficient model for $R_{\text{cta}}$. As for extremely low-resource languages where the Judge model completely fails to comprehend the input, the CTA reward would indeed be unreliable. In such cases, we suggest falling back to only Format+Acc+LC reward or replacing it with rule-based CTA rewards. Considering our paper mainly focuses on language consistency and the effectiveness of the CTA reward on multilingual reasoning, we leave the study of extremely low-resource languages for further study.
@@ Appendix A Instruction for Cross-lingual Thinking Alignment Reward
# Task Analyze and quantify the consistency of key intermediate results between an English and a [target] thought process for a given math problem. # Inputs I will provide you with three items: [English Math Problem]: The original problem in English. [English Thought Process]: The step-by-step reasoning for solving the problem in English. [[target] Thought Process]: The step-by-step reasoning for solving the problem in [target]. # Instructions You must perform the following analysis internally: Identify all key intermediate results from the [English Thought Process]. Key results include variable definitions, equations, critical calculation values, and the final answer. For each key result identified in English, find its mathematical equivalent in the [[target] Thought Process]. Calculate the consistency score using the following formula: Score = (Number of matched, mathematically equivalent pairs) / (Total number of key results identified in the English process)
# Output Format Your final output MUST BE a single decimal number between 0 and 1. And the number should be wrapped by <score> and </score>. Do NOT include any text, explanation, titles, analysis, or any other characters. The response must only be the number itself wrapped by <score> and </score>. Example of a valid response: <score>0.925</score> ——- [English Math Problem]: [en-question] [English Thought Process]: [en-think] [[target] Thought Process]: [x-think]
@@ B.4 Instructions of No-Training Baselines
Appendix C Additional Results
In-Domain Languages Out-of-Domain Languages
Methods ja ko fr pt th ID-AVG en es ar vi zh OOD-AVG ALL-AVG
Metric: Language Consistency (LC, %)
Qwen3-4B-Thinking-2507 0.00 0.00 0.00 0.00 0.00 0.00 99.79 0.00 0.00 0.00 65.52 33.06 16.53
Prompt-Control (No Training) 0.00 0.00 0.00 0.00 0.00 0.00 99.94 0.00 0.02 0.02 68.50 33.70 16.85
DIT (No Training) 95.13 16.66 2.37 94.31 9.44 43.58 99.94 97.71 64.13 98.72 82.61 88.62 66.10
QRT (No Training) 99.13 97.17 96.87 92.35 96.31 96.37 99.96 99.32 96.83 99.35 84.21 95.94 96.15
Cold-Start SFT 68.73 54.72 16.64 3.97 17.64 32.34 89.50 9.89 1.66 24.18 64.67 37.98 35.16
Naive-RL 0.00 0.00 0.00 0.00 0.00 0.00 96.01 0.00 0.00 0.00 10.87 21.38 10.69
SLC-RL 76.00 65.02 78.94 71.72 80.71 74.48 87.99 67.70 61.75 69.34 68.20 71.00 72.74
M-Thinker-4B $\Rightarrow$ Iter-1 (Ours) 99.61 99.86 99.46 99.44 99.90 99.66 99.42 98.53 99.61 99.33 82.52 95.88 97.77
Metric: Accuracy (Acc, %)
Qwen3-4B-Thinking-2507 76.15 77.30 78.70 79.17 76.06 77.48 84.36 80.15 74.51 77.81 75.42 78.45 77.96
Prompt-Control (No Training) 75.72 76.85 78.10 79.70 75.37 77.15 85.10 77.95 73.66 77.35 74.30 77.67 77.41
DIT (No Training) 70.63 72.07 77.88 77.20 75.44 74.64 80.55 78.72 70.74 73.01 78.20 76.24 75.44
QRT (No Training) 69.16 67.14 75.78 78.14 66.80 71.40 78.12 80.65 66.30 74.80 76.86 75.34 73.37
Cold-Start SFT 64.41 68.72 78.55 80.24 75.77 73.54 83.05 77.95 76.09 77.23 76.90 78.24 75.89
Naive-RL 76.07 77.13 77.57 78.29 75.82 76.98 77.52 78.11 76.32 79.03 77.60 77.72 77.35
SLC-RL 58.78 58.47 71.82 75.60 60.89 65.11 81.11 75.37 52.83 68.25 75.51 70.61 67.86
M-Thinker-4B $\Rightarrow$ Iter-1 (Ours) 71.41 68.11 76.38 74.64 67.23 71.56 82.24 78.63 71.23 73.27 78.57 76.79 74.17
Metric: Language Consistency & Accuracy (LC&Acc, %)
Qwen3-4B-Thinking-2507 0.00 0.00 0.00 0.00 0.00 0.00 84.36 0.00 0.00 0.00 47.40 26.35 13.18
Prompt-Control (No Training) 0.00 0.00 0.00 0.00 0.00 0.00 85.03 0.00 0.02 0.02 49.59 26.93 13.47
DIT (No Training) 66.67 10.31 2.37 72.97 8.84 32.23 80.51 76.86 44.89 72.35 63.33 67.59 49.91
QRT (No Training) 68.96 65.32 73.68 71.99 64.08 68.81 78.10 80.39 65.73 74.59 63.91 72.54 70.68
Cold-Start SFT 50.88 41.79 12.94 3.20 13.12 24.39 82.99 8.50 1.62 17.40 50.51 32.21 28.30
Naive-RL 0.00 0.00 0.00 0.00 0.00 0.00 74.20 0.00 0.00 0.00 10.46 16.93 8.47
SLC-RL 56.25 48.44 66.37 62.33 59.68 58.61 81.07 59.93 36.46 51.86 52.14 56.29 57.45
M-Thinker-4B $\Rightarrow$ Iter-1 (Ours) 71.03 67.96 75.84 74.11 67.17 71.22 81.87 77.51 70.85 73.23 65.36 73.76 72.49
Table 8: The LC, Acc, and LC&Acc (%) results on the MMATH benchmark of the Qwen3-4B-Thinking-2507 backbone. “ID-avg/OOD-avg” is the average result of five In-Domain/Out-of-Domain languages and “ALL-AVG” is the average result of all ten languages. The result in bold means the best result.
In-Domain Languages Out-of-Domain Languages
Methods ja ko fr pt th ID-AVG en es ar vi zh OOD-AVG ALL-AVG
Metric: Language Consistency (LC, %)
DIT (No Training) 15.34 0.29 48.41 44.85 3.54 22.48 90.25 32.91 4.18 27.64 61.50 43.30 32.89
QRT (No Training) 12.21 0.08 52.72 41.60 10.71 23.47 90.97 33.39 9.34 26.41 67.19 45.46 34.46
Metric: Accuracy (Acc, %)
DIT (No Training) 19.39 17.41 31.51 28.66 18.53 23.10 47.52 27.10 11.49 17.31 39.09 28.50 25.80
QRT (No Training) 14.89 16.51 28.16 30.06 6.68 19.26 45.55 26.10 10.25 16.67 42.01 28.11 23.69
Metric: Language Consistency & Accuracy (LC&Acc, %)
DIT (No Training) 7.83 0.06 25.99 23.36 0.55 11.56 47.10 22.38 1.93 13.74 33.23 23.68 17.62
QRT (No Training) 6.10 0.06 25.22 25.17 1.86 11.68 45.45 21.61 2.48 13.66 36.17 23.87 17.78
LC Acc LC&Acc
Settings ID-AVG OOD-AVG ALL-AVG ID-AVG OOD-AVG ALL-AVG ID-AVG OOD-AVG ALL-AVG
w/ Hard-LC (1.5B) 99.16 92.44 95.80 31.72 39.85 35.78 31.68 37.18 34.43
w/ Soft-LC (1.5B) 0.00 37.16 18.58 46.80 49.16 47.98 0.00 19.47 9.74
w/ Hard-LC (7B) 99.34 92.27 95.80 55.56 62.09 58.83 55.47 58.16 56.81
w/ Soft-LC (7B) 76.00 74.24 75.12 56.64 60.55 58.59 43.14 46.88 45.01
Table 10: The detailed comparison between the "Hard-LC" ( with LC/Format/Acc rewards) and "Soft-LC" (with SLC/Format/Acc rewards).
Format Language Consistency Accuracy SLC-RL SLC-RL-s Ours (w/o $R_{\text{cta}}$) Ours
/ / 0 -1 -1 -1
$\checkmark$ 1*(0+0)=0 1*(-1+0)=-1 -1 -1
$\checkmark$ $\checkmark$ 1*(0+0.9)=0.9 1*(-1+1)=0 -1 -1
[…]
Table 11: The detailed reward scores of different methods on different scenarios. “SLC-RL-s” means the same scale of soft-LC reward with our $R_{\text{lc}}$. “Ours” represents M-thinker-1.5B (Iter-1).
LC Acc LC&Acc
Methods LC/Acc Reward ID-AVG OOD-AVG ALL-AVG ID-AVG OOD-AVG ALL-AVG ID-AVG OOD-AVG ALL-AVG
SLC-RL [0,0.1]+[0,0.9] 0.00 37.16 18.58 46.80 49.16 47.98 0.00 19.47 9.74
SLC-RL-s [-1,0]+[0,1] 39.34 57.46 48.40 44.66 48.78 46.72 16.95 28.96 22.96
Ours (w/o $R_{\text{cta}}$) [-1,0]&[0,1] 99.16 92.44 95.80 31.72 39.85 35.78 31.68 37.18 34.43
Ours [-1,0]&[0,1]($R_{\text{cta}}$) 99.19 84.48 91.83 35.59 44.22 39.90 35.39 38.37 36.88
Table 12: The results of SLC-RL with different reward magnitudes. “SLC-RL-s” means the same scale of soft-LC reward with our $R_{\text{lc}}$. “Ours” represents M-thinker-1.5B (Iter-1).
@@ Appendix C Additional Results
M-Thinker-7B $\Rightarrow$ Iter-2 (Ours) 98.75 95.30 99.65 99.00 94.65 97.47 97.55 98.65 64.80 100.00 89.25 90.05 93.76
@@
DeepSeek-R1-Distill-Qwen-1.5B 13.60 15.77 18.62 18.73 11.25 15.59 21.23 19.47 14.36 18.25 20.00 18.66 17.13
M-Thinker-1.5B $\Rightarrow$ Iter-1 (Ours) 16.64 12.33 23.03 23.40 12.77 17.63 30.13 23.23 15.90 17.42 25.08 22.35 19.99
M-Thinker-1.5B $\Rightarrow$ Iter-2 (Ours) 19.65 17.39 24.87 24.76 16.84 20.70 32.41 25.63 19.11 20.83 27.98 25.19 22.95
DeepSeek-R1-Distill-Qwen-7B 28.45 31.72 35.41 33.12 27.72 31.28 36.93 33.51 28.93 31.96 30.67 32.40 31.84
M-Thinker-7B $\Rightarrow$ Iter-1 (Ours) 29.99 28.59 35.02 35.30 29.35 31.65 40.80 34.72 29.83 31.99 34.80 34.43 33.04
M-Thinker-7B $\Rightarrow$ Iter-2 (Ours) 35.24 33.92 40.02 39.73 34.40 36.66 42.48 38.33 37.08 37.19 42.73 39.56 38.11
@@
DeepSeek-R1-Distill-Qwen-1.5B 0.67 0.00 3.03 3.36 0.16 1.44 21.15 3.71 0.57 2.40 16.09 8.78 5.11
M-Thinker-1.5B $\Rightarrow$ Iter-1 (Ours) 16.43 12.07 23.03 23.27 12.75 17.51 29.89 23.18 4.55 16.96 22.87 19.49 18.50
M-Thinker-1.5B $\Rightarrow$ Iter-2 (Ours) 19.61 17.04 24.86 24.51 16.80 20.56 32.36 25.44 1.55 20.59 24.62 20.91 20.74
DeepSeek-R1-Distill-Qwen-7B 3.05 1.89 5.89 4.44 2.39 3.53 36.70 5.74 2.45 4.89 25.27 15.01 9.27
M-Thinker-7B $\Rightarrow$ Iter-1 (Ours) 29.81 28.55 34.97 35.11 29.11 31.51 40.67 34.61 27.18 31.97 31.73 33.23 32.37
M-Thinker-7B $\Rightarrow$ Iter-2 (Ours) 35.13 33.13 39.96 39.36 33.79 36.27 42.48 37.96 24.18 37.19 39.24 36.21 36.24
@@
C.1 Results of Qwen3-4B-Thinking-2507
We list the LC, Acc, and LC&Acc (%) results on the MMATH benchmark of the Qwen3-4B-Thinking-2507 backbone in Table 8. The results demonstrate the superiority of our method on both language consistency and answer accuracy over other baselines. Due to the time and resource limitations, we only conduct GRPO training for one iteration. The effectiveness of the iterative training strategy has been proven in the other backbones.
C.2 Detailed Results of DeepSeek-R1-Distill-Qwen-1.5B
We report the LC, Acc, and LC&Acc (%) results on the MMATH benchmark of the DeepSeek-R1-Distill-Qwen-1.5B backbone for each language in Table 9.
C.3 Hard vs. Soft Language Consistency Reward
To compare the performance of hard/soft LC reward, we conduct GRPO training from the same cold-start SFT model and report the results in Table 10. The results reveal distinct behaviors across model sizes. For the 1.5B Model, the hard LC constraint ensures higher language consistency but leads to a drop in answer accuracy (Acc) compared to the soft LC reward, as the smaller model struggles to satisfy strict language constraints while reasoning correctly. For the 7B Model, the accuracy gap becomes negligible (the ALL-AVG “58.83” of Hard-LC even surpasses the one “58.59” of Soft-LC). Additionally, the Hard-LC reward significantly outperforms the Soft-LC reward in maintaining language consistency (“95.8” vs. “75.12”). In summary, the performance achieved with either hard-LC or soft-LC ultimately depends on the model’s inherent capability: the stronger the model, the better hard-LC can simultaneously ensure linguistic consistency and answer accuracy.
C.4 SLC-RL with Different Reward Magnitudes
For a clear comparison, we also conduct SLC-RL with the same reward magnitude as ours. First, we list the detailed reward scores of different methods in Table 11. Specifically, “SLC-RL-s” utilizes the same LC-reward magnitude as our method; however, it still underperforms our method as shown in Table 12. The results demonstrate that the hard LC reward facilitates higher language consistency, and the effectiveness of our M-Thinker stems from the strict LC reward and the CTA reward rather than merely the LC reward scale.
C.5 Results of PolyMath
Appendix D Additional Analysis
D.1 Detailed Ablation Results
@@
D.2 Other Alternative Judge Metrics of the CTA Reward
MAPO She et al. (2024) utilizes the NLLB model to calculate translation probabilities between English and multilingual responses, serving as a selection criterion to construct preference data for DPO training. While this translation-consistency approach can be adapted as a reward function for cross-lingual alignment, we suspect that it may not be an effective additional reward signal from two primary concerns: (1) Translation models often degrade when processing long chain-of-thought (CoT) reasoning heavily interspersed with mathematical formulas (LaTeX). This may result in erroneous reward signals and lead to unstable RL training. As evidenced in Table 4 of CM-Align Zhang et al. (2025b), MAPO struggles to maintain high reward accuracy across different models. (2) Fundamentally, a translation-based reward encourages the model to generate literal translations of the English reasoning path. This forces the model into "translationese" rather than allowing it to “think natively” in the target language. In contrast, our CTA reward checks for the logical equivalence of key intermediate steps. This allows the model to employ native reasoning patterns and syntactic structures as long as the key steps remain correct.
@@ D.2 Other Alternative Judge Metrics of the CTA Reward
D.3 Annotation Guidelines of Human Evaluation
The annotation follows a rigorous two-step procedure:
  • Reference Extraction: First, experts proficient in English extract key intermediate reasoning steps from the reference English thinking paths.
  • Alignment Verification: Experts proficient in the target languages are then provided with the non-English thinking paths and the extracted English key steps. They independently identify key intermediate steps in the target language and calculate the alignment ratio with the English counterparts.